### v5

# Scaling LLM Planning: NL2FLow for Parametric Problem Generation and Rigorous Evaluation

Abstract
Robust workflow composition is critical for effective agent performance, yet progress in Large Language Model (LLM) planning and reasoning is hindered by a scarcity of scalable evaluation data. This work introduces NL2Flow, a fully automated pipeline for generating and evaluating workflow planning problems. NL2Flow generates problems parametrically in a structured intermediate representation, translating them into both natural language and formal PDDL. I evaluate several open-source, instruct-tuned LLMs on a dataset of 2296 low-difficulty problems generated by NL2Flow. Results demonstrate that the best-performing model achieved 86% success in generating valid plans and 69% in generating optimal plans (for solvable problems). Regression analysis shows that the influence of problem characteristics on plan generation is contingent on both model and prompt design. Importantly, translating natural language problems into a structured JSON representation prior to symbolic planning significantly improved success rates, suggesting a benefit from neuro-symbolic integration. These findings underscore the importance of understanding error sources within LLM reasoning as systems scale to more complex tasks. As LLM reasoning scales to increasingly complex problems, understanding the shifting bottlenecks and sources of error within these systems will be crucial.
1 Introduction
Agentic systems, increasingly deployed to automate complex business processes, rely on the ability to dynamically compose tools and agents into effective workflows. The core challenge in building these systems is planning: determining the optimal sequence of actions to achieve a given goal. While traditional AI planning techniques offer formal guarantees, they struggle with the scalability and adaptability demanded by real-world applications involving a large number of tools and rapidly changing environments.
Figure 1: Steps to obtain a plan from a planning problem in natural language. Two approaches are shown: (1) Direct planning, where a Large Language Model (LLM) Planner directly generates a plan from the input. (2) Intermediate representation planning, where an LLM Translator first converts the natural language instruction into an intermediate representation, which is then used by symbolic planning algorithms to compute a plan.
Traditional AI planning systems rely on formal representations, such as those defined by Planning Domain Definition Language (PDDL). While these systems can offer guarantees like optimality, building them requires painstaking knowledge engineering to translate natural language goals into formal problem descriptions (georgievski2015htn, ). This translation process is a known bottleneck, and the complexity of maintaining and scaling these systems increases substantially with the number of tools and task complexities. For example, business automation systems can involve thousands of APIs. Consequently, symbolic planners struggle to find optimal workflows within reasonable timeframes, often requiring developers to manually reduce the scope of the problem by limiting the number of APIs considered.
Large Language Models (LLMs) present a promising alternative, demonstrating impressive capabilities in natural language understanding, generation, and reasoning. However, using LLMs for reliable planning remains challenging. Existing approaches often resort to free-form plan generation, which lacks logical consistency, or rely on reinforcement learning confined to limited environments (planningopen, ). A key advancement would be the ability to reliably translate natural language planning problems into formal representations or to directly generate logically sound and executable plans (Figure 1). Successful LLM-based planning would significantly reduce the engineering effort associated with building robust natural language planning systems.
Despite their potential, evaluating and improving LLM planning capabilities is hampered by the scarcity of scalable data generation and evaluation methodologies. This study addresses this need by introducing NL2Flow (https://github.com/IBM/nl2flow), a scalable data generation and evaluation method for workflows. NL2Flow uses random and parametric approaches to create a dataset of workflow problems, eliminating the need for human or LLM annotation and enabling the creation of large-scale datasets. I demonstrate that NL2Flow can generate workflow problems in a domain where open-sourced, instructed-tuned LLMs without task-specific optimization or architectural modifications exhibit promising planning performance for low-difficulty problems. Establishing a domain in which large language models (LLMs) exhibit promising reasoning and planning capabilities without task-specific optimization and modifications can enable researchers to investigate the causes of their domain-dependent reasoning/planning performance.
Importantly, the use of PDDL allows for symbolic evaluation of generated plans, providing mathematically grounded metrics, which include optimality, for reliable and scalable assessment. By systematically varying goals and agent/tool configurations, NL2Flow facilitates evaluation of LLMs both as planners and as translators from natural language to workflow representations, highlighting key bottlenecks in building robust, natural language-grounded AI planning systems (Figure 1). In this study, I focus on applying LLMs to low-difficulty workflow generation problems, which are readily solvable by symbolic planners, as a first step towards tackling more complex scenarios.
2 Method
This study used an automated planning integration system, NL2Flow, to generate a dataset of workflow planning problems and evaluate the performance of large language models in generating workflows and translating workflow problems in natural language. The NL2Flow’s profiler module (profiler.generators.batch_data_generator) was used to programmatically generate a diverse set of planning scenarios, varying parameters. This allowed for controlled generation of data suitable for evaluating LLM performance across a range of workflow configurations. LLM inference was conducted on compute clusters equipped with NVIDIA H100, H200, and A100 GPUs (with larger models, for example, DeepSeek-V3, using up to 8 H200 GPUs with 400GB memory). LLM-generated workflows were evaluated using NL2Flow’s Debugger module (nl2flow.debug.debug.BasicDebugger), which assesses soundness, validity, and optimality. A description of the AI planning principles is provided in the Appendix.
2.1 Domain Description
Listing 1: NL2Flow domain example for booking a business trip (https://github.com/IBM/nl2flow)
[0] ask(conference name)
[1] ask(username)
[2] invoice = Registration Bot(conference name, username)
[3] map(invoice, conference registration)
[4] name, address, Employee ID, Passport = Travel Info Agent(username)
[5] visa = Visa Application(Passport, address)
[6] ask(start date)
[7] ask(end date)
[8] map(home, address)
[9] map(BOS, destination)
[10] booking = Taxi(date, address, destination)
[11] map(LAX, address)
[12] map(JW Marriott Los Angeles LA 900 W Olympic Blvd, destination)
[13] booking = Taxi(date, address, destination)
[14] map(end date, date)
[15] map(JW Marriott Los Angeles LA 900 W Olympic Blvd, address)
[16] map(LAX, destination)
[17] booking = Taxi(date, address, destination)
[18] map(BOS, address)
[19] map(destination, home)
[20] booking = Taxi(date, address, destination)
[21] flight_ticket, hotel_booking = Concur(start date, end date, home, destination)
[22] map(flight_ticket, ticket to conference)
[23] assert \$hotel_booking.price + \$flight_ticket.price < 1500
[24] approval = Trip Approval(ticket to conference, conference registration)
Automated workflow completion was a core requirement for the intelligent systems investigated in this study. I focused on the NL2Flow domain, which combined service composition with goal-oriented dialogue to generate task workflows (chakraborti2022natural, ). The objective was to develop a system capable of autonomously selecting and composing a sequence of actions, which are termed ‘custom’ actions, representing available services, to achieve user goals (e.g., submitting an expense report) (Listing 1). Each ‘custom‘ action was modeled as a parameterized function, and successful execution necessitated providing values for a defined set of input parameters. Critically, these parameter values were frequently unknown at the initiation of a workflow, representing information the system needed to acquire from the user. To address this, the system used a dialogue-based approach. When a parameter value was missing, the system generated an ‘ask’ action, a dialogue act specifically designed to elicit the necessary information. This process of identifying and requesting missing parameter values was analogous to slot-filling, a common technique in task-oriented dialogue systems, and ensured the system possessed the information required to progress. Successfully executing a ‘custom‘ action updated the system’s internal state, transitioning the corresponding parameter from unknown to known, and enabling the system to proceed with subsequent actions in the workflow.
2.2 Parametric Data Generation
A parametric workflow problem generator was developed to create an intermediate representation (IR) of the problem space. Implemented in Python, this generator allows for controlled variation in workflow complexity and characteristics. A dataset of 2,296 workflows was generated using a fixed random seed (20250429) to ensure reproducibility. Workflow generation was governed by six key parameters (Table 1): the number of ‘custom‘ actions (representing available operations), ‘custom‘ action arity (the number of input/output parameters defining each action’s interface), the number of goal actions (specifying the desired outcome), the proportion of slot-fillable variables (variables requiring values during planning), the complexity of variable mapping, and the proportion of coupled actions (actions where outputs serve as inputs to subsequent actions, creating dependencies). I aimed to generate ten unique samples for each valid parameter combination. ‘custom‘ actions were defined as callable components with specified input and output parameter sets, invoked when all required input parameters are known, and producing corresponding output parameters. These actions represent operations commonly found in agentic frameworks, such as agents and tools.
Table 1: Workflow data generator configuration
Parameters Values
‘custom‘ actions [2, 3, 4, 5]
Arity [1, 2, 3]
Coupling [0, 0.375, 0.626, 1.0]
Slot-fillable variables [0.375, 0.626, 1.0]
Mapping [0]
Goals [1, 2]
To construct a workflow problem, 30 variables were generated and designated as either slot-fillable or initially known. ‘custom‘ actions were then created and connected to establish dependencies, with action outputs serving as inputs for subsequent actions. Variables not directly used within any action were assigned pre-defined values. A ‘custom‘ action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the ‘custom‘ action set. To focus on core workflow generation challenges, we excluded variable mapping and initial variable values from this study. The resulting intermediate representation (Python object) was compiled into PDDL strings using NL2Flow, enabling symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication.
The parametric workflow problem generator constructed an intermediate representation (Listing A1 in Appendix), implemented as a Python object, to define the problem space. This representation comprises five key components:
  • available_agentsavailable\_agents represent the available ‘custom‘ actions
  • actuator_signatureactuator\_signature specifies the input and output parameters for each agent.
  • mappingsmappings enable the transfer of values between variables.
  • available_dataavailable\_data define the known values of variables at the workflow’s outset.
  • goal_agent_idsgoal\_agent\_ids are the ids of the desired actions to be achieved.
From this intermediate representation, I generated natural language prompts and a formal PDDL representation of the workflow problem. See Appendix, Prompt Generation and PDDL Compilation for details on prompt generation and PDDL compilation, respectively. Language model responses were parsed according to the procedure outlined in Appendix, Language Model Response Parsing.
2.3 Plan Evaluation
NL2Flow evaluated plans generated by LLMs using the Planning Domain Definition Language (PDDL). LLM-generated plans were extracted from text and converted to PDDL via NL2Flow, using an intermediate representation from the data generator. Details of this PDDL compilation process are provided in the PDDL Compilation subsection of the Appendix. The resulting PDDL formulations were then assessed with a symbolic planner. For clarity, a sound plan is defined as one that is executable; a valid plan is executable and achieves the goal state; and an optimal plan is executable, reaches the goal state, and does so with minimal cost. When a planning problem lacks a feasible solution, any proposed plan is invalid. Furthermore, identifying an optimal plan requires the existence of at least one feasible solution.
To assess the soundness of generated plans, the pddl compilation omits goal achievement. NL2Flow’s approach distinguishes between soundness and validity. For soundness checking, the planning problem was compiled into PDDL, but instead of including the original goals as requirements, the actions reported as executed within the LLM’s response were incorporated. This approach tests the LLM’s claimed actions actually executable. Validity was checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan. In this low-difficulty workflow problem setting, a plan was invalid, also, if the system prompted the user for a non-slot-fillable variable, or if a ‘custom‘ action was invoked with unknown input parameter values. I investigated the use of VAL (howey2004val, ) for plan validation; however, VAL did not support the PDDL language features necessary to represent the complexities introduced by NL2Flow’s PDDL compilation process, precluding its use in this evaluation.
[…]
Following normalization of the explanatory variables (Tables LABEL:tab:lrnoplanlong, LABEL:tab:lrnoplanshort, LABEL:tab:lrplanvaliditylong, LABEL:tab:lrplanvalidityshort, LABEL:tab:lrplanoptimallong, and LABEL:tab:lrplanoptimalshort in Appendix), I used logistic regression, implemented in the Python statsmodels package, to assess the contribution of these variables to different aspects of LLM workflow generation capacity.
3 Results
3.1 Planning Task
3.1.1 Plan Length
[…]
Kolmogorov-Smirnov tests indicated that prompt style did not significantly alter plan length distributions for DeepSeek-V3 and Llama-3.1-405b-instruct-fp8, but did significantly affect them for other models (Figures A1 and A2 in Appendix). Mann-Whitney U tests further showed that prompt style did not significantly influence the median plan length for DeepSeek-V3, Llama-3.1-405b-instruct-fp8, and Llama-3.3-70b-instruct, while it did impact the median plan length for the remaining models.
3.1.2 Sound, Valid, and Optimal Planning
Overall, Llama-3.3-70B-instruct, when paired with the verbose prompt style, demonstrated the strongest planning performance (Figure 2). This model achieved the highest rates of generating sound plans (80%), valid plans (77%), and optimal plans (64%) among all tested models. Using the concise prompt style resulted in a noticeable decrease in performance, with rates dropping to 71%, 60%, and 46% respectively. A similar, though more pronounced, trend was observed for Llama-3.1-8B-instruct.
Figure 2: Comparison of LLM performance in generating sound, valid, and optimal plans using verbose and concise prompt styles
3.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 3) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid plans was largely attributable to its ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of ‘custom‘ actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – showed that factors influencing the LLM’s failure to correctly identify ’no plan’ scenarios varied depending on the model and the prompt style (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort in Appendix). However, no single variable consistently predicted failure across all models or prompt styles.
Figure 3: Comparison of large language model performance (accuracy) in detecting ’no plan’ scenarios, evaluated with verbose and concise prompt styles
3.1.4 Feasible Plan
When restricted to problems with feasible plans, Llama-3.1-405B-instruct-fp8 demonstrated the strongest performance, achieving the high rates of generating sound (91%), valid (85%), and optimal plans (69%) using the concise prompt style (Figure 4). I further investigated the influence of problem characteristics – including optimal plan length, number of ‘custom‘ actions, arity, coupling, proportion of slot-fillable variables, and number of goals – on the generation of valid plans. The relationship between these explanatory variables and plan quality (soundness, validity, and optimality) varied considerably across language models.
Figure 4: Performance comparison of large language models on planning tasks with feasible plans
Across most models, optimal plan length was the strongest predictor of valid plan generation (Tables LABEL:tab:lrplanvaliditylong and LABEL:tab:lrplanvalidityshort in Appendix). However, this relationship was not consistent. With the verbose prompt style, logistic regression with L1 regularization ($\alpha=1$) failed to identify any significant predictors of valid plan generation for Granite-3.3-8B-instruct and Llama-3.1-8B-instruct. Likewise, when using the concise prompt, optimal plan length was not a significant predictor for Llama-3.3-70B-instruct, Mixtral-8x22B-instruct-v0.1, and Qwen2.5-72B-instruct. These results indicate that the influence of problem difficulty on valid workflow generation is model-dependent.
Regarding optimal plan generation, coupling consistently explained a significant portion of the variance across all models when the verbose prompt was used (Table LABEL:tab:lrplanoptimallong in Appendix). Meanwhile, this trend was observed only in DeekSeek-V3 and Llama-3.3-70B-instruct when using the concise prompt style (Table LABEL:tab:lrplanoptimalshort in Appendix). These results suggest that the impact of problem characteristics on optimal plan generation is both model- and prompt-dependent.
3.2 Natural Language to JSON Translation
DeekSeek-V3 and Llama-3.1-405B-instruct-FP8 consistently generated valid JSON strings (100%) (Figure 5). Specifically, all JSON outputs from these models included all required elements and contained no hallucinated information. In contrast, DeepSeek-coder-33B-instruct produced valid JSON reliably, exhibiting a mean of over 6 hallucinated elements and over 6 missing elements. The models Granite-3.3-8B-instruct, Llama-3.1-8B-instruct, and Qwen2.5-72B-instruct failed to generate valid JSON in over 90% of attempts.
4 Related Work
Current evaluation of reasoning capacity often focuses on final answer accuracy for benchmarks like MATH500 (lightman2023let, ), AIME24, GSM8K (cobbe2021training, ), and MATH. However, this metric overlooks the validity of the reasoning ’process’ itself; a correct answer can be achieved through flawed reasoning, effectively rewarding a successful, yet invalid, plan. Furthermore, relying solely on final answer correctness leaves the quality and structure of the reasoning process opaque. Similarly, BIRD (li2023can, ), a natural language to SQL translation benchmark, primarily relies on final answer accuracy, despite the inherent planning aspects of SQL query construction alongside elements like intent detection. GSM-Symbolic (mirzadeh2024gsm, ) attempts to address this by generating mathematical problems from symbolic templates, allowing for manual verification of solutions. ReasonAgain (yu2024reasonagain, ) uses GPT-4o to generate Python programs for GSM8K and MATH, validating reasoning by executing the generated code.
Beyond these benchmarks, several studies use “puzzle” datasets (shojaee2025illusion, ), which include Tower of Hanoi, river crossing, blocks world, and checker jumping, to assess LLM reasoning. These puzzles offer the advantage of verifiable reasoning steps through simulation and can be scaled to challenge LLMs as they saturate. However, the applicability of these abstract puzzles to real-world scenarios remains limited.
A growing body of work focuses on evaluating and training LLMs for workflow automation and reasoning. Several datasets have been proposed to assess LLM planning capabilities, differing in their generation methods (static vs. dynamically generated) and evaluation approaches (manual vs. programmatic). This section details relevant work, highlighting these key distinctions.
Figure 5: LLM-based natural language to JSON translation for planning
4.1 Static Datasets & Primarily Manual/LLM-as-Judge Evaluation
FlowBench (xiao2024flowbench, ) presents a static dataset of 51 scenarios across 6 domains, focusing on multi-turn interactions. Evaluation relied heavily on human verification by comparing expected and observed traces, identifying a need for improved LLM planning. Similarly, HuggingGPT (shen2023hugginggpt, ) uses a static, human-annotated dataset augmented with GPT-4 generated pseudo-labels, and uses LLM-as-a-judge for graph tasks. TaskLAMA (yuan2024tasklama, ) also provides a static dataset, created through human annotation, for structured complex task decomposition. ISG-BENCH (chen2024interleaved, ) focuses on interleaved text and image generation, using a static dataset and relying on MLLM-as-a-judge for scoring alignment, coherence, and helpfulness, grounded in human-annotated truth. WikiPlan and RecipePlan contribute static text-image pairs for task completion. These datasets, while valuable, often require significant human effort for creation and/or evaluation (lu2023multimodal, ).
4.2 Static Datasets & Programmatic Evaluation
TaskBench (shen2024taskbench, ) divides task automation into decomposition, tool selection, and parameter prediction, leveraging a static dataset of 17,331 samples generated with GPT-4. The authors used programmatic metrics (F1, Normalized Edit Distance, Rouge) for evaluation, with some partial manual inspection. This approach offers scalability due to a large sample size but can be limited by the quality and coverage of the initial dataset. Open Grounded Planning (planningopen, ) also aims to generate executable plans, providing LLMs with an extensive action library and leveraging existing datasets. Evaluation combines programmatic metrics, such as executability and pass rate, with LLM-as-a-judge for assessing completeness, feasibility, and relevance.
4.3 Algorithmically Generated Datasets & Programmatic Evaluation
WORFBENCH (qiao2024benchmarking, ) represents a step towards automated evaluation through its data generator for graph workflow structures. LLM responses are evaluated via subgraph matching. However, there is a limitation: the need to define the entire graph upfront, hindering scalability – a limitation that search-based methods, like my own, avoid by not requiring complete graph definition for data generation and evaluation.
My work extends existing research by presenting a fully automated pipeline for data generation and evaluation, overcoming the limitations of static datasets and enabling scalable, programmatic assessment of LLM workflow generation. However, the effectiveness of this approach is ultimately bounded by the capacity of the symbolic planner to evaluate workflow optimality.
5 Conclusion
This work demonstrates that current open-source, instruct-tuned LLMs can generate effective plans directly from natural language instructions, outperforming previous LLM planning studies that relied on formal PDDL representations (silver2022pddl, ; valmeekam2023planning, ; kokel2025acpbench, ). This difference likely arises from the nature of the reasoning tasks: our study focused on workflow creation, where generated actions inherently involve fixed input and output variables, contrasting with the flexible variable binding common in domains like the Blocks World. Importantly, the complexity of workflow creation is comparable to that of domains used in prior LLM planning research. The observed pattern reinforces the hypothesis that LLM planning capabilities are domain-specific, potentially rooted in their statistical reliance on token prediction and the associated sensitivity to the compositional complexity of input tokens (zhao2025chain, ).
Regression analysis confirmed that the influence of problem characteristics on the goodness of plan is dependent on both the language model and the prompting strategy employed. This highlights the importance of model- and prompt-specific analysis, cautioning against broad generalizations across LLMs. My study demonstrated that translating natural language into an intermediate JSON workflow representation significantly improved performance compared to direct plan generation (Figure 1). This suggests that prematurely replacing established neuro-symbolic planning systems with direct language-to-action reasoning may lead to a decline in plan quality. While decomposing planning into multiple steps can introduce complexity, simplifying the system by directly reasoning from language to action can, in some applications, prioritize reduced complexity over plan quality. However, as LLM reasoning capabilities advance and scale to more complex problems, the limitations and error sources within AI planning systems are likely to evolve.
NL2Flow extends beyond simple LLM assessment by providing a framework for evaluating the planning capabilities of novel agentic architectures. These architectures vary in agent and tool configurations and operate with randomly assigned goals, expressed as LLM-generated utterances based on system-defined agent and tool descriptions (including those adhering to the Model Context Protocol (MCP)). By translating these goals and configurations into PDDL, NL2Flow enables the evaluation of agentic system trajectories using a standardized planning framework, allowing for assessment of an agent’s ability to select and compose tools based on custom, symbolically-verifiable workflows. Given the prevalence of chain-of-thought, tree-of-thought, and self-consistency prompting methods in agentic systems for workflow creation, verifying the soundness of this semantic approach is crucial for improving system reliability.
Future work will focus on expanding this automated pipeline to generate datasets exhibiting greater combinatorial complexity and incorporating diverse constraints. Investigating the interplay between prompting strategies, model architecture, and dataset characteristics will be crucial for understanding how to best use LLMs for complex workflow automation. Ultimately, a systematic analysis of LLM limitations across a spectrum of reasoning tasks will be vital for guiding future development and realizing the full potential of these models.
Appendix
Intermediate Representation Example
Listing A1: Intermediate representation from data generator. An example of the intermediate representation is provided in the supplementary material as ‘intermediate_representation.json‘.
"available_agents": [
"agent_id": "a__2",
"actuator_signature": {
"in_sig_full": [
"name": "v__7",
"data_type": null,
"mappable": false,
"slot_fillable": true
"out_sig_full": [
"name": "v__9",
"data_type": null,
"mappable": false,
"slot_fillable": true
"goal_agent_ids": [
"a__1",
"a__0"
"mappings": [],
"available_data": [
"v__10",
null
Prompt Generation
This study examined the capacity of LLMs to generate executable plans to achieve specified goals and to translate natural language problem descriptions into JSON schemas. I adopted a few-shot prompting approach, constructing prompt styles for both tasks and providing six example demonstrations. Crucially, each example within the prompt presented a natural language description of the plan execution before the corresponding plan expressed in the target format (Listing A2). This was intended to guide the LLM by demonstrating the logical connection between action sequences and their executable representations, thereby increasing the likelihood of generating valid and goal-directed plans.
Listing A2: Planning response example
PLAN EXPLANATION
0. Execute action a__2. This will result in acquiring v__1. v__1 is required later by action A.
1. Execute action a__1 with $a$ as input. This will result in acquiring v__2. v__2 is required later by action a__3. Since executing A was a goal of this plan, return the results of A(v__1) to the user.
2. Execute action a__3 with v__2 as input. Since executing a__3 was a goal of this plan, return the results of a__3(v__2) to the user.
PLAN
[0] v__1 = a__2()
[1] v__2 = a__1(v__1)
[2] a__3(v__2)
For the planning task, I investigated the impact of prompt style, comparing ‘verbose‘ (Listing A5) and ‘concise‘ (Listing A6) prompt styles. The ‘verbose‘ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ‘concise‘ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. Listings A3 and A4 show the description of ‘custom‘ action a__1a\_\_1 in verbose and concise prompts, respectively. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt.
Listing A3: ‘custom‘ action in verbose prompt
Action a__1 has input parameters v__11, v__12, and v__13 and it outputs variables v__15, v__16, and v__18. To execute action a__1, variables v__11, v__12, and v__13 must be known. After executing action a__1, variables v__15, v__16, and v__18 are known.
Listing A4: ‘custom‘ action in concise prompt
name: a__1
inputs: v__11, v__12, v__13
outputs: v__15, v__16, v__18
For translation, a single prompt style (Listing A7) was used, designed to elicit accurate and fluent translations. The complete prompt styles used for each task – verbose_planning_prompt.txt, concise_planning_prompt.txt, and translation_prompt.txt – are included in the supplementary material to facilitate reproducibility, enable independent verification of my results, and encourage further research in this area. Each prompt style followed a consistent structure comprising three sections: Context, Instruction, and Examples (Listings A5 and A6). The Context section provided a programmatically generated natural language description of the problem. The Instruction section detailed the desired planning process, specifying preferences for plan structure, the required output format, and guidance for handling unsolvable problems. To facilitate parsing and execution, plans were to be expressed in a Python-like syntax (e.g., outputs = ⟨action name⟩(⟨parameters⟩)). I standardized action names as a__⟨number⟩and variable names as v__⟨number⟩to promote consistency and reduce ambiguity.
Listing A5: Verbose planning prompt
CONTEXT
The system has actions a__0, a__1, and a__2.
Action a__1 has input parameters v__13, v__12, and v__11 and it outputs variables v__15, v__16, and v__18. To execute action a__1, variables v__13, v__12, and v__11 must be known. After executing action a__1, variables v__15, v__16, and v__18 are known.
Action a__2 has
INSTRUCTION
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as ’noplan’
Here are some examples.
Example #1
Now output your plan explanation and plan.
{plan_explanation_and_plan}
Listing A6: Concise planning prompt
CONTEXT
The system has actions a__0, a__1, and a__2. To execute an action the values of their inputs must be known.
actions:
name: a__1
inputs: v__11, v__12, v__13
outputs: v__15, v__18, v__16
name: a__2
inputs:
INSTRUCTION
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as ’noplan’
Here are some examples.
Example #1
Now output your plan explanation and plan.
{plan_explanation_and_plan}
This study evaluated nine LLMs across translation and planning tasks: CodeLlamma-34B-instruct, DeepSeek-coder-33B-instruct, DeepSeek-V3(685B, fp8), Granite-3.3-8B-instruct, Llama-3.1-8B-Instruct, Llama-3.1-405B-instruct-fp8, Llama-3.3-70B-instruct, Mixtral-8x22B-instruct-v0.1, and Qwen2.5-72B-instruct.
Listing A7: Translation prompt
CONTEXT
The system has actions a__0, a__1, and a__2.
Action a__1 has input parameters v__13, v__12, and v__11 and it outputs variables v__15, v__16, and v__18. To execute action a__1, variables v__13, v__12, and v__11 must be known. After executing action a__1, variables v__15, v__16, and v__18 are known.
Action a__2 has
INSTRUCTION
Write a json to describe the context.
Here are some examples.
Example #1
Now output your json.
{json}
Language Model Response Parsing
To establish a lower bound on large language model performance, I used minimal parsing techniques for both translation and planning tasks, prioritizing simplicity and avoiding potentially biasing post-processing steps.
For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. The initial intermediate representation schema used nested structures to represent ‘custom‘ actions; however, this complexity proved unnecessary, as the same information could be represented with a simplified schema. To evaluate the capacity of LLMs to translate natural language descriptions of workflow problems into JSON, I used a streamlined schema (Listing A8) defined as follows:
  • actionsactions represents ‘custom‘ actions.
  • mappingsmappings corresponds to the ’mappings’ present in the intermediate representation.
  • available_dataavailable\_data is a list of variables with known values at the workflow’s initiation.
  • askable_parametersaskable\_parameters is a list of slot-fillable variables for user input.
  • unaskable_parametersunaskable\_parameters is a list of variables for which the system is prohibited from prompting the user.
  • goal_action_idsgoal\_action\_ids corresponds to the ’goal_action_ids’ in the intermediate representation.
An example of this schema, used for evaluating LLM translation capacity, is provided as ‘translation_object.json‘ in the supplementary material.
For the planning task, each action within a generated plan was extracted and stored as a string within a Python list. This list was subsequently used as input to the NL2Flow debugger, a symbolic planning validation tool that assesses the logical consistency and feasibility of the generated plan.
Listing A8: JSON schema used for evaluating LLM’s translation capacity
"actions": [
"id": "a__1",
"input": [
"v__1",
"v__2"
"output": [
"v__3"
"mappings": [],
"available_data": [
"v__5",
"askable_parameters": [
"v__1",
"unaskable_parameters": [
"v__4",
"goal_action_ids": [
"a__3"
AI Planning
AI planning addresses the problem of finding a sequence of actions that transforms an initial state into a goal state. Formally, a planning problem is defined as $P=\langle P,F,D,I,G\rangle$, where:
  • $P$ is a finite set of predicates describing properties of the world.
  • $F$ is a finite set of numeric variables.
  • $D$ is the domain model, comprising a set of action schemas.
  • II is the initial state, describing the world at the start of planning.
  • $G$ is the goal state, specifying the desired outcome.
I adopt the closed-world assumption, meaning any fact not explicitly stated in the initial state is considered false. An action $a$ is defined by its preconditions and effects: $a=\langle pre(a),eff(a)\rangle$. Preconditions, $pre(a)$, define the conditions that must hold in a state for the action to be applicable. Effects, $eff(a)$, specify how the state changes upon execution of the action.
A state $s$ is a conjunction of predicates and numeric variable assignments: $s=\langle P_{s},V_{s}\rangle$, where $P_{s}\subseteq P$ represents the set of true predicates, and $V_{s}\in\mathbb{Q}^{|F|}$ represents the assignment of rational values from $\mathbb{Q}$ to each numeric variable in $F$.
Conditions are expressed as a combination of predicate literals and numeric constraints. A numeric constraint is a relation between arithmetic expressions: $\langle exp_{1},\triangleright,exp_{2}\rangle$, where ${<,,>,,=}\triangleright\in\{$$\leq$$\geq,$=\}. Arithmetic expressions (expexp) are built from numeric variables in $F$, rational constants in $\mathbb{Q}$, and the operators +, and -.
An effect $eff(a)$ is a triple $\langle P^{+}_{eff},P^{-}_{eff},U_{eff}\rangle$, where:
  • $P^{+}_{eff}\subseteq P$ is the set of predicates that become true.
  • $P^{-}_{eff}\subseteq P$ is the set of predicates that become false.
  • $U_{eff}$ is a set of numeric updates, each of the form $\langle f,:=/+=/-=,exp\rangle$, where $f\in F$ is the updated variable and expexp is the new value or change in value.
Applying an action $a$ in state $s$ is possible if all predicates in $pre(a)$ are present in $P_{s}$ and all numeric constraints in $pre(a)$ hold true when evaluated using $V_{s}$. The resulting state $s^{\prime}=\langle P^{\prime}_{s},V^{\prime}_{s}\rangle$, is computed as follows: $P^{\prime}_{s}=(P_{s}\setminus P^{-}_{eff})\cup P^{+}_{eff}$, and each variable $f\in F$ is updated in $V^{\prime}_{s}$ according to the operator and expression specified in $U_{eff}$.
A plan $\pi$ is a sequence of actions. The trajectory of a plan, starting from initial state $I$, is the sequence of states $s_{0},s_{1},\dots$
PDDL Compilation
This study used NL2Flow to translate an intermediate representation of a workflow creation problem into PDDL. The PDDL formulation within NL2Flow enabled cost-based planning to identify efficient workflows (plans). While NL2Flow was designed to represent a broad range of workflow automation concepts, this study focused on generating PDDL for simple, low-difficulty workflows. For instance, the mapping capacity feature, which allowed variable values to be shared, was excluded from the current implementation.
Planning problems in this work were formalized using the Planning Domain Definition Language (PDDL). PDDL separated each problem’s description into a domain model, which defined the possible actions and their effects, and a problem instance, specifying the initial state and goal conditions. Example domain and problem files used in this study (domain.pddl and problem.pddl, respectively) were provided as supplementary material. A domain model included declarations of predicates, functions, and actions, while a problem instance defined the initial state through a set of grounded predicates and function assignments. Any fact not explicitly stated as true in the initial state was considered false. Grounding involved replacing variables in predicates and functions with concrete values.
Types
The domain defines several types:
  • genericgeneric: Represents the objects being processed – the entities whose slots need filling.
  • operatoroperator: Represents actions the agent can take.
  • hasdonestatehas-done-state: Represents temporal states (past, present, future) to track whether an action has been performed.
  • constraintstatusconstraint-status: Represents certainty levels (True, False) likely used in constraints or conditions.
  • datumstatedatum-state: Represents the certainty of information about an object (certain, uncertain, unknown).
  • labellabel: (Currently unused) labeling objects or slots.
  • objectobject: A base type for extensibility.
Constants
Many constants are defined for each type. The ’generic’ constants (v__xv\_\_x) represent specific instances of the objects being processed. The hasdonestatehas{-}done{-}state constants track the temporal status of actions. The constraintstatusconstraint{-}status and datumstatedatum{-}state constants define the possible values for these attributes. operatoroperator constants (a__xa\_\_x) represent specific, pre-defined actions.
Predicates
The predicates define the relationships and properties within the domain:
  • (has_done?xoperator?x2hasdonestate)(has\_done\>?x\>-\>operator\>?x2\>-\>has{-}done{-}state): Indicates that an operator has been performed in a specific time state.
  • (been_used?xgeneric)(been\_used\>?x\>-\>generic): Marks an object as having been processed.
  • (new_item?xgeneric)(new\_item\>?x\>-\>generic): Indicates an object is newly introduced.
  • (known?xgeneric?x2datumstate)(known\>?x\>-\>generic\>?x2\>-\>datum{-}state): Represents the agent’s knowledge about an object’s certainty. This is central to the planning process.
  • (not_slotfillable?xgeneric)(not\_slotfillable\>?x\>-\>generic): Indicates an object cannot have its slots filled (perhaps due to inherent properties).
  • (is_mappable?xgeneric?x2generic)(is\_mappable\>?x\>-\>generic\>?x2\>-\>generic) and (not_mappable?xgeneric?x2generic)(not\_mappable\>?x\>-\>generic\>?x2\>-\>generic): The ability to map one object to another.
  • (mapped?xgeneric)(mapped\>?x\>-\>generic): Indicates an object has been mapped.
  • (not_usable?xgeneric)(not\_usable\>?x\>-\>generic): Indicates an object cannot be used.
  • (mapped_to?xgeneric?x2generic)(mapped\_to\>?x\>-\>generic\>?x2\>-generic): Indicates an object is mapped to another object.
  • (free?xgeneric)(free\>?x\>-\>generic): Indicates an object is available for processing.
  • (done_goal_pre)(done\_goal\_pre) and (done_goal_post)(done\_goal\_post): Flags to indicate the goal has been achieved before and after certain actions.
Functions
Functions are used to represent numeric state variables. They allow you to model quantities that change over time and are used in numeric effects of actions.
  • (totalcost)(total{-}cost): A numeric function representing the cumulative cost of the plan.
  • (slot_goodness?xgeneric)(slot\_goodness\>?x\>-\>generic): A numeric function that likely evaluates the q̈ualityör usefulness of filling a slot for a given object.
  • (affinity?xgeneric?x2generic)(affinity\>?x\>-\>generic\>?x2\>-\>generic): A numeric function that likely measures the similarity or relatedness between two objects.
Actions
The domain defines two action types. ‘custom‘ actions are named according to the pattern a__xa\_\_x.
  • a__xa\_\_x:
    • Precondition: Require that the action hasn’t been done in the past and that specific objects (v__xv\_\_x) are known with certainty.
    • Effect: Mark the action as done in the present. Mark the object as used. Free up objects. Mark the newly freed object as known with certainty. Mark the newly freed object as not mapped. Increase the totalcosttotal{-}cost by 10.
Listing A9: ‘custom‘ action example in PDDL
(:action a__2
:parameters ()
:precondition (and (not (has_done a__2 past)) (known v__7 certain))
:effect (and
(has_done a__2 present)
(been_used v__7)
(free v__9)
(known v__9 certain)
(not (mapped v__9))
(increase (total-cost ) 10)
For example, Listing A9 illustrate the structure of ‘custom‘ actions defined in the domain model. This action, a__2a\_\_2, has no parameters. Its precondition requires that the action hasn’t been performed in the past (not(has_donea__2past)not(has\_done\>a\_\_2\>past)) and that a variable v__7v\_\_7 is known with certainty (knownv__7certainknown\>v\_\_7\>certain). The action’s effect marks the action as done in the present (has_donea__2presenthas\_done\>a\_\_2\>present), indicates v__7v\_\_7 has been used, makes v__9v\_\_9 free, establishes v__9v\_\_9 as known with certainty, indicates v__9v\_\_9 is no longer mapped, and increases the $total\mbox{-}cost$ by 10.
The askask action prompts the system to request information from the user if parameters are slot-fillable (Listing A10).
Listing A10: ‘ask‘ action example in PDDL
(:action ask
:parameters (?x - generic)
:precondition (and (not (known ?x certain)) (not (not_slotfillable ?x))
:effect (and
(free ?x)
(mapped_to ?x ?x)
(known ?x certain)
(not (not_usable ?x))
(not (mapped ?x))
(increase (total-cost ) (slot_goodness ?x))
  • ‘ask‘:
    • Takes a genericgeneric object as a parameter.
    • Precondition: Requires that the object is not currently known with certainty and is slotfillable (notnot_slotfillablenot\>not\_slotfillable).
    • Effect: Frees the object. Maps the object to itself (a placeholder). Marks the object as known with certainty. Marks the object as usable (not not_usablenot\_usable). Marks the object as not mapped. Increases the totalcosttotal-cost by the slot_goodnessslot\_goodness of the object.
Problem
A problem describes a specific instance of a planning problem. It works in conjunction with a domain, which defines the general rules and actions applicable to a class of problems. A problem contains:
  • Initial state: The starting situation for the planner. This is a description of what’s true at the beginning. Predicates not explicitly defined as true in the initial state of a problem are considered false (closed-world assumption).
  • Goal state: What the planner needs to achieve to solve the problem. This is a condition that must be true when the plan is complete.
  • Specifies objects: The specific objects that exist in this particular problem instance.
The initial condition of the problem, as described in ‘problem.pddl‘ in supplementary material, can be summarized as follows:
  • slot_goodnessslot\_goodness is defined for all generic objects (v__xv\_\_x and new_object_generic_0new\_object\_generic\_0).
  • totalcosttotal-cost is initialized to 0.0.
  • For any generic object v__xv\_\_x used as an input parameter to a ‘custom‘ action, the predicate $(been_{u}sed\>v\_\_x)$ is true.
  • The predicate (new_itemnew_object_generic_0)(new\_item\>new\_object\_generic\_0) is true.
  • For all generic object, predicates (mapped_to?x?x)(mapped\_to\>?x\>?x) are true. In other words, All generic objects are self-mapped.
In this study, the initial state was defined by a set of variables (v__xv\_\_x), each assigned a slot_goodnessslot\_goodness value of 1000000 if slot-fillable, and 150000 if not. These differing slot_goodnessslot\_goodness values represented a ’soft’ constraint, which was not used in this study because the precondition for the ‘ask‘ action prevented the system from requesting information for variables that were not slot-fillable. The mapped_tomapped\_to predicate establishes a self-mapping for each resource, essentially stating that each resource is initially assigned to itself. A new_itemnew\_item predicate indicates the creation of a new resource. The goal is to achieve conditions to complete a task or tasks (and(has_donea__xpresent)(has_donea__xpresent)and\>(has\_done\>a\_\_x\>present)(has\_done\>a\_\_x\>present)) while minimizing the totalcosttotal-cost.
Regression Analysis Results
Table A1: Logistic regression analysis for no plan with verbose prompt style
DeepSeek-V3
Pseudo R-squ Log-Likelihood
-0.1845 -48.236
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -20.9808 14.949 -1.403 0.16 -50.281 8.319
arity -19.1153 15.123 -1.264 0.206 -48.755 10.524
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -2.3799 11.24 -0.212 0.832 -24.409 19.649
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
-0.0137 -146.9
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -0.0534 10.769 -0.005 0.996 -21.161 21.054
arity -15.7939 11.914 -1.326 0.185 -39.144 7.556
coupling -6.1441 4.01 -1.532 0.126 -14.004 1.716
slot-fillable variables -8.00E-17 11.144 -7.18E-18 1 -21.841 21.841
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
-0.03309 -110.76
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -14.7038 12.09 -1.216 0.224 -38.4 8.993
arity -11.7652 13.046 -0.902 0.367 -37.335 13.805
coupling 0 nan nan nan nan nan
slot-fillable variables -2.3115 13.035 -0.177 0.859 -27.86 23.237
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
0.007233 -189.65
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 11.4596 8.745 1.31 0.19 -5.681 28.6
arity -7.33E-17 8.774 -8.36E-18 1 -17.196 17.196
coupling -2.0446 3.507 -0.583 0.56 -8.918 4.829
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
-0.08881 -65.393
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -21.3022 13.77 -1.547 0.122 -48.291 5.686
arity -17.1386 13.92 -1.231 0.218 -44.421 10.144
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -8.56E-17 10.337 -8.28E-18 1 -20.261 20.261
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
-0.05837 -57.104
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -37.7334 5.594 -6.746 0 -48.697 -26.77
arity 0 nan nan nan nan nan
coupling -3.1613 5.535 -0.571 0.568 -14.009 7.687
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.01343 -92.919
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -0.7135 12.438 -0.057 0.954 -25.091 23.664
arity -24.2876 12.676 -1.916 0.055 -49.132 0.557
coupling -6.8198 4.877 -1.398 0.162 -16.378 2.739
slot-fillable variables 0 nan nan nan nan nan
goals -1.7047 9.061 -0.188 0.851 -19.464 16.054
Table A2: Logistic regression analysis for no plan with concise prompt style
DeepSeek-V3
Pseudo R-squ Log-Likelihood
inf -18.971
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -27.0837 17.651 -1.534 0.125 -61.679 7.512
arity -22.9006 17.806 -1.286 0.198 -57.800 11.999
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -2.53e-16 13.455 -1.88e-17 1.000 -26.372 26.372
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.0352 -99.331
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -4.953 4.509 -1.099 0.272 -13.79 3.884
slot-fillable variables -10.9061 7.956 -1.371 0.17 -26.5 4.688
goals -16.5726 7.71 -2.15 0.032 -31.684 -1.461
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
-0.02809 -61.746
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -4.3217 15.636 -0.276 0.782 -34.968 26.325
arity -1.1087 16.917 -0.066 0.948 -34.266 32.049
coupling -1.6357 5.5 -0.297 0.766 -12.416 9.144
slot-fillable variables -15.4646 16.811 -0.92 0.358 -48.414 17.485
goals -17.7946 10.376 -1.715 0.086 -38.131 2.542
Llama-3-3-70B-instruct
Pseudo R-squ Log-Likelihood
0.005184 -194.65
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -4.4632 3.174 -1.406 0.16 -10.684 1.757
slot-fillable variables 0 nan nan nan nan nan
goals -5.4489 3.137 -1.737 0.082 -11.597 0.699
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
-0.03562 -106.62
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -3.6785 11.656 -0.316 0.752 -26.524 19.167
arity -19.0184 11.842 -1.606 0.108 -42.228 4.191
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -7.0985 8.556 -0.83 0.407 -23.868 9.671
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
0.09344 -91.359
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -7.5232 4.342 -1.733 0.083 -16.033 0.986
slot-fillable variables 0 nan nan nan nan nan
goals -27.1984 4.446 -6.118 0 -35.912 -18.485
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.1284 -64.37
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -18.1791 14.71 -1.236 0.217 -47.01 10.652
arity -14.1435 15.761 -0.897 0.37 -45.034 16.747
coupling 0 nan nan nan nan nan
slot-fillable variables -6.4428 16.611 -0.388 0.698 -39 26.114
goals 0 nan nan nan nan nan
Table A3: Logistic regression analysis for validity of plans for problems with feasible plans with verbose prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.08473 -1043.2
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -41.7807 5.085 -8.216 0 -51.748 -31.814
‘custom‘ actions 62.4231 6.135 10.174 0 50.398 74.448
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 17.53 6.108 2.87 0.004 5.558 29.502
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
-0.0004765 -1375.9
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.1782 -767.54
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -64.5546 5.771 -11.186 0 -75.866 -53.243
‘custom‘ actions 95.0417 7.206 13.19 0 80.919 109.165
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 30.7633 6.989 4.402 0 17.066 44.461
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
0.09968 -931.42
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -49.7385 5.343 -9.309 0 -60.211 -39.266
‘custom‘ actions 63.1362 6.384 9.89 0 50.624 75.649
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 35.2596 6.511 5.416 0 22.499 48.02
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.003312 -1371.4
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -3.4977 2.824 -1.238 0.216 -9.034 2.038
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0.375 2.823 0.133 0.894 -5.157 5.907
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
0.02272 -1062
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -23.8761 5.058 -4.72 0 -33.79 -13.962
‘custom‘ actions 50.4787 6.101 8.274 0 38.521 62.437
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 16.9506 6.157 2.753 0.006 4.883 29.018
goals 0 nan nan nan nan nan
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
0.03135 -1023.5
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -22.2533 5.146 -4.324 0 -32.34 -12.167
‘custom‘ actions 56.6528 6.277 9.025 0 44.349 68.956
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 13.7726 6.279 2.193 0.028 1.465 26.08
goals 0 nan nan nan nan nan
Table A4: Logistic regression analysis for validity of plans for problems with feasible plans with concise prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.1384 -723.47
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -49.3192 5.858 -8.419 0.000 -60.801 -37.837
‘custom‘ actions 94.7476 7.746 12.232 0.000 79.566 109.929
arity 0 nan nan nan nan nan
coupling 3.997e-16 3.945 1.01e-16 1.000 -7.732 7.732
slot-fillable variables 26.3420 7.395 3.562 0.000 11.848 40.836
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.01306 -1352.2
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -5.658 2.837 -1.994 0.046 -11.219 -0.097
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -5.5151 2.842 -1.941 0.052 -11.084 0.054
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.1559 -713.1
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -56.0808 5.887 -9.527 0 -67.618 -44.543
‘custom‘ actions 99.9576 7.79 12.832 0 84.69 115.225
arity 0 nan nan nan nan nan
coupling 1.2189 3.931 0.31 0.757 -6.486 8.924
slot-fillable variables 25.7175 7.369 3.49 0 11.275 40.16
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
-0.01675 -1300
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 9.1485 3.029 3.02 0.003 3.211 15.086
slot-fillable variables 8.7886 3.01 2.92 0.003 2.89 14.687
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.08154 -339.31
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -4.00E-16 17.191 -2.32E-17 1 -33.694 33.694
‘custom‘ actions -32.2708 9.824 -3.285 0.001 -51.525 -13.017
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -99.841 20.12 -4.962 0 -139.276 -60.406
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
-0.008253 -1185.3
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 29.6337 4.892 6.057 0 20.045 39.223
goals 8.0143 4.786 1.674 0.094 -1.366 17.395
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.03739 -834.25
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -3.29E-15 5.972 -5.51E-16 1 -11.705 11.705
‘custom‘ actions 50.0114 7.328 6.825 0 35.648 64.374
arity 0 nan nan nan nan nan
coupling 4.4832 4.009 1.118 0.263 -3.374 12.341
slot-fillable variables 19.8769 7.324 2.714 0.007 5.522 34.232
goals 0 nan nan nan nan nan
Table A5: Logistic regression analysis for optimality of plans for problems with feasible plans with verbose prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.02228 -1342.1
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -1.8931 4.346 -0.436 0.663 -10.411 6.625
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -16.0413 3.061 -5.24 0 -22.041 -10.041
slot-fillable variables 7.9159 4.59 1.725 0.085 -1.08 16.912
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.04275 -1236.8
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions -2.7211 3.102 -0.877 0.38 -8.801 3.359
arity 0 nan nan nan nan nan
coupling -28.8463 3.278 -8.799 0 -35.272 -22.421
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.05037 -1236.7
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -6.761 4.742 -1.426 0.154 -16.055 2.533
‘custom‘ actions 22.0291 5.675 3.882 0 10.906 33.152
arity 0 nan nan nan nan nan
coupling -22.9507 3.236 -7.093 0 -29.292 -16.609
slot-fillable variables 21.0509 5.726 3.676 0 9.828 32.274
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
0.07216 -1209.5
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -25.7043 6.722 -3.824 0 -38.88 -12.529
‘custom‘ actions 11.1087 5.649 1.966 0.049 0.037 22.181
arity 18.4486 6.866 2.687 0.007 4.992 31.905
coupling -18.039 3.248 -5.554 0 -24.405 -11.673
slot-fillable variables 28.5103 5.866 4.86 0 17.013 40.007
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.0603 -1023.4
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions -28.2565 3.368 -8.39 0 -34.858 -21.655
arity 0 nan nan nan nan nan
coupling -30.1313 3.653 -8.248 0 -37.291 -22.971
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
0.04701 -1302.3
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -0.3057 6.459 -0.047 0.962 -12.965 12.354
‘custom‘ actions 0 nan nan nan nan nan
arity 17.5519 6.677 2.629 0.009 4.465 30.638
coupling -21.6725 3.125 -6.935 0 -27.798 -15.547
slot-fillable variables 5.1429 4.821 1.067 0.286 -4.306 14.592
goals 0 nan nan nan nan nan
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
0.05514 -1299.6
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -24.8377 3.085 -8.051 0 -30.884 -18.791
slot-fillable variables 9.6276 3.015 3.194 0.001 3.719 15.536
goals 0 nan nan nan nan nan
Table A6: Logistic regression analysis for optimality of plans for problems with feasible plans with concise prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.009851 -1279.7
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 20.4360 4.515 4.527 0.000 11.588 29.284
arity 4.1622 4.290 0.970 0.332 -4.246 12.570
coupling -22.6719 3.180 -7.130 0.000 -28.904 -16.440
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.01306 -1352.2
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -5.658 2.837 -1.994 0.046 -11.219 -0.097
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -5.5151 2.842 -1.941 0.052 -11.084 0.054
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.1559 -713.1
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -56.0808 5.887 -9.527 0 -67.618 -44.543
‘custom‘ actions 99.9576 7.79 12.832 0 84.69 115.225
arity 0 nan nan nan nan nan
coupling 1.2189 3.931 0.31 0.757 -6.486 8.924
slot-fillable variables 25.7175 7.369 3.49 0 11.275 40.16
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
-0.01675 -1300
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 9.1485 3.029 3.02 0.003 3.211 15.086
slot-fillable variables 8.7886 3.01 2.92 0.003 2.89 14.687
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.08154 -339.31
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -4.00E-16 17.191 -2.32E-17 1 -33.694 33.694
‘custom‘ actions -32.2708 9.824 -3.285 0.001 -51.525 -13.017
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -99.841 20.12 -4.962 0 -139.276 -60.406
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
-0.008253 -1185.3
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 29.6337 4.892 6.057 0 20.045 39.223
goals 8.0143 4.786 1.674 0.094 -1.366 17.395
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.03739 -834.25
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -3.29E-15 5.972 -5.51E-16 1 -11.705 11.705
‘custom‘ actions 50.0114 7.328 6.825 0 35.648 64.374
arity 0 nan nan nan nan nan
coupling 4.4832 4.009 1.118 0.263 -3.374 12.341
slot-fillable variables 19.8769 7.324 2.714 0.007 5.522 34.232
goals 0 nan nan nan nan nan
LLM Plan Length Distributions
Figure A1: Distribution of LLM-generated plan lengths for different prompt styles
[…]